# Top Score on the Wrong Exam: On Benchmarking in Machine Learning for Vulnerability Detection

### v1

@@ Abstract
According to our survey of the machine learning for vulnerability detection (ML4VD) literature published in the top Software Engineering conferences, every paper in the past 5 years defines ML4VD as a binary classification problem:
@@
In this paper, we ask whether this decision can really be made without further context and study both vulnerable and non-vulnerable functions in the most popular ML4VD datasets. A function is vulnerable if it was involved in a patch of an actual security flaw and confirmed to cause the vulnerability. It is non-vulnerable otherwise. We find that in almost all cases this decision cannot be made without further context. Vulnerable functions are often vulnerable only because a corresponding vulnerability-inducing calling context exists while non-vulnerable functions would often be vulnerable if a corresponding context existed.
But why do ML4VD techniques perform so well even though there is demonstrably not enough information in these samples? Spurious correlations: We find that high accuracy can be achieved even when only word counts are available. This shows that these datasets can be exploited to achieve high accuracy without actually detecting any security vulnerabilities.
We conclude that the current problem statement of ML4VD is ill-defined and call into question the internal validity of this growing body of work. Constructively, we call for more effective benchmarking methodologies to evaluate the true capabilities of ML4VD, propose alternative problem statements, and examine broader implications for the evaluation of machine learning and programming analysis research.
@@ I Introduction
In our survey of all 22 ML4VD papers published at the Top-4 Software Engineering conferences over the last five years [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], we find that state-of-the-art ML4VD techniques exclusively define ML4VD as a binary classification problem: Given an isolated function $f$, decide whether $f$ contains a security vulnerability. The technique with the lowest classification error on the evaluation dataset is considered the best at detecting security vulnerabilities.
@@ 21 TF_LITE_ENSURE_STATUS(context->ResizeTensor(context, output, output_dims));
However, based on our experience, we hypothesized that it might not always be possible to determine whether a function is vulnerable or not without additional context. We call these vulnerabilities context-dependent. Consider the example in Figure 1. If this function from the DiverseVul benchmark dataset [23] is called with num_splits set to zero, it will crash with a division-by-zero in Line 7. However, without knowing whether this function can ever be called with num_splits set to zero, we cannot reliably decide if the division-by-zero could actually be observed. This function parameter might as well be properly validated by every caller function. This context-dependency problem is well-known in static analysis [24, 25], and software testing [26].
In this paper, we set out to quantify the prevalence of context-dependent vulnerabilities in the most popular datasets and end up revealing a fundamental flaw in the most-widely used evaluation methodology that is underpinning the progress of the nascent research area of ML4VD. We find that the vulnerability of a function cannot be decided without further context for more than 90% of functions. This includes functions with both types of labels, vulnerable or secure: If the right context existed, the function would be considered vulnerable. Respectively, only because the right context exists, the function is considered vulnerable.
Given our findings, we conclude that the current problem statement of ML4VD as a function-level classification problem is inadequate. The reported results in the literature, which are based on this problem statement, may not accurately reflect the true capabilities of the evaluated techniques at the task of vulnerability detection. In other words, there is currently no evidence that ML4VD techniques are actually capable of identifying security vulnerabilities in functions.
But why do ML4VD techniques perform well at this binary classification task when there is demonstrably not enough information in over 90% of samples (even after addressing label inaccuracies)? We identify spuriously correlated features as a potential reason. Training simple models, like a gradient boosting classifier using only word counts and disregarding code structure, we achieved results comparable to those of state-of-the-art ML4VD models. This suggests that ML4VD techniques only appear to perform well due to the chosen evaluation methodology. During classification, ML4VD techniques rely on spuriously correlated features to achieve high scores and do not genuinely detect vulnerabilities.
@@
    We analyze all papers published at the Top-4 Software Engineering conferences over the last five years and find that state-of-the-art ML4VD techniques exclusively define ML4VD as a function-level classification problem.
@@
    Why do ML4VD techniques still perform well at the function-level classification problem? We demonstrate that they may rely on spurious features to achieve high scores without genuinely detecting vulnerabilities.
  • $\star$
    We publish all of our code and results for reproducibility. They are available at https://github.com/niklasrisse/TopScoreWrongExam.
@@ II Background
The problem that the vulnerability of a function may depend on context that is external to the given function (i.e., the context-dependency problem) is well known in static analysis and software testing.
@@
Static Analysis. We distinguish between inter-procedural and intra-procedural analysis, where the former is concerned with the analysis of the entire system and the latter with the analysis of individual functions. Intra-procedural static analysis tools often struggle with false positives due to context-dependency. For example, a static analyzer might flag potential bugs based on the analysis of code patterns but cannot always discern the specific conditions under which a bug manifests [27, 28]. Hence, Le et al. [24] propose to differentiate between manifest bugs, which are context-independent, and latent bugs, which depend on preconditions in the calling context. Manifest bugs can be reported without making strong assumptions about the calling context, whereas latent bugs cannot. To address these challenges, tools like PhASAR [25] leverage an inter-procedural analysis to detect security vulnerabilities that depend on external context.
@@
Software Testing. We distinguish between system-level testing and unit-level testing, where the former is concerned with test cases for the entire system and the latter with test cases for individual system units, like functions. Automated unit test generation often falls short of identifying issues that only occur under specific conditions or in particular environments. As noted by Harrold and Orso [29], unit tests can produce false positives when they are not adequately designed to account for the broader context in which a function operates. This can lead to an inflated number of reported issues that are not actual bugs, thereby complicating the debugging process. Hence, property-based testing [30, 31, 32] requires users to define function preconditions in addition to assertions to ensure that the assertions fire only under valid function parameters (i.e., under valid context). To summarize, both the static analysis and software testing research communities acknowledge the problem of context-dependency and address it through various techniques aimed at reducing the number of false positives.
Machine Learning (ML4VD). Sejfia et al. [4] realize that a focus on individual base units (e.g., functions) prevents us from learning to find vulnerabilities that emerge from the interaction of multiple base units (MBU). Given only vulnerabilities that required multiple functions to be fixed (called MBU vulnerabilities), they observe a drastic drop in detection accuracy of the evaluated ML4VD models. Croft et al. [33] study the quality of benchmark datasets in the ML4VD literature and, as part of their manual re-labeling policy, decide that a function is conservatively also considered as vulnerable if it invokes a function that is known to be vulnerable–which raises an interesting context-dependency question for us. They find that a large percentage of functions in the most widely used benchmarks are incorrectly labeled as vulnerable. In this paper, we reproduce their experiment to identify and study actually vulnerable functions in terms of their context-dependence. We study the prevalence of ML4VD as a binary classification problem separately in the next section.
However, to the best of our knowledge, no work exists that studies whether the vulnerability of the base units in the available datasets can be decided without further context in the first place. Addressing this gap, our paper aims to uncover the shortcomings of current evaluation methodologies and emphasizes the urgent need for more context-aware approaches to accurately assess the true capabilities of ML4VD techniques in detecting security vulnerabilities.
Spurious features. Arp et al. [34] provide initial evidence that machine learning techniques can learn to predict labels correctly based on artifacts in code snippets without addressing the actual security task at hand (spurious correlation). Similarly, Risse et al. [35] demonstrate that ML4VD techniques overfit to label-unrelated features by utilizing semantic preserving transformations of code. Building on this, our work further explores the extent to which these techniques depend on spurious features, suggesting that the reliance on such features could provide an alternative explanation for the high accuracy scores reported in the literature.
@@ III Literature Survey
Figure 2: Number of papers published at the top software engineering conferences over the last five years (until July 2024) that present or evaluate ML4VD techniques.
@@ III-A Methodology
To ensure a thorough and unbiased analysis of the recent literature on machine learning for vulnerability detection (ML4VD) in the field of software engineering, we adopted a systematic methodology based on established guidelines for conducting literature reviews in software engineering research [36, 37]. Our approach encompassed three primary phases: defining the scope, selecting relevant papers, and systematically analyzing the selected papers.
@@
Scope. Our literature survey focuses on the past five years, targeting publications from 2020 to 2024. Given the prominence and impact of certain venues in the field of software engineering, we selected the four top-tier conferences—International Conference on Software Engineering (ICSE), Foundations of Software Engineering (FSE), International Symposium on Software Testing and Analysis (ISSTA), and Automated Software Engineering (ASE). These conferences were chosen due to their reputation for publishing high-quality and influential research.
Criteria for Paper Selection. To identify relevant papers, we established clear selection criteria. The papers must either propose a novel ML4VD technique or evaluate existing ML4VD techniques. To facilitate the identification process, we defined a set of keywords that are indicative of the research focus in this area: vulnerability, vulnerable, detect, detection, discovery, machine, learning, artificial, intelligence, AI, ML, deep, graph, neural, network, large, language, model.
Process. We conducted a systematic search of the accepted papers lists from the targeted conferences. For each conference, we utilized the Safari browser search function to scan the titles of all accepted papers for the presence of our predefined keywords. Any title containing one of the keywords was flagged for further review.
In the next step, we filtered the flagged titles to specifically identify those that propose or evaluate machine learning techniques for vulnerability detection (ML4VD). We manually checked each of these titles and, where necessary, reviewed the abstracts and full texts to assess their relevance based on our selection criteria.
@@
To answer research questions 1) and 2), we documented how ML4VD is defined, and the datasets used for empirical evaluation within each identified ML4VD paper.
@@ III-B Results
Our literature survey identified 22 papers that met our selection criteria and were included in our analysis [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Note that accepted papers for ASE 2024 have not been released at the submission time of this paper, which is why they are not included in the results.
Figure 3: Datasets that are used by ML4VD papers published at the top software engineering conferences over the last five years (until July 2024).
The distribution of these papers across the years and conferences, as displayed in Figure 2, indicates a clear trend: the number of papers focused on ML4VD is increasing annually. This upward trajectory underscores the growing interest and significance of this research area within the software engineering community. Additionally, we found a large variety of machine learning techniques being employed, utilizing different data representations (graph-based, token-based), model architectures (e.g., large language models, graph neural networks, convolutional neural networks), and learning algorithms (e.g., contrastive learning).
Problem Statement. To our surprise, all 22 papers define ML4VD as a binary function-level classification problem. This problem statement, which focuses on determining whether a given function contains a security vulnerability, has become the de facto standard in the field.
Datasets. Regarding datasets, the majority of the papers relied on a limited set of popular datasets for empirical evaluation. Figure 3 shows all datasets that were used by the 22 papers and the number of times they were used. Specifically, BigVul [38] was used by 16 papers, Devign [39] was used by 14 papers, and ReVeal [40] by 10 papers. Notably, 21 out of the 22 papers utilized at least one of these three datasets, reflecting their dominance in the field.
Literature Survey. All 22 papers analyzed over the past five years from top software engineering conferences (ICSE, FSE, ISSTA, and ASE) define ML4VD as a binary classification problem: Given an isolated function $f$, decide whether $f$ contains a security vulnerability. Moreover, three datasets—BigVul, Devign, and ReVeal—are overwhelmingly favored, with 21 out of the 22 papers using at least one of these datasets for empirical evaluation.

### v2

@@ Abstract
According to our survey of machine learning for vulnerability detection (ML4VD), 9 in every 10 papers published in the past five years define ML4VD as a function-level binary classification problem:
@@
From our experience as security researchers, faced with deciding whether a given function makes the program vulnerable to attacks, we would often first want to understand the context in which this function is called.
In this paper, we study how often this decision can really be made without further context and study both vulnerable and non-vulnerable functions in the most popular ML4VD datasets. We call a function “vulnerable” if it was involved in a patch of an actual security flaw and confirmed to cause the program’s vulnerability. It is “non-vulnerable” otherwise. We find that in almost all cases this decision cannot be made without further context. Vulnerable functions are often vulnerable only because a corresponding vulnerability-inducing calling context exists while non-vulnerable functions would often be vulnerable if a corresponding context existed.
But why do ML4VD techniques achieve high scores even though there is demonstrably not enough information in these samples? Spurious correlations: We find that high scores can be achieved even when only word counts are available. This shows that these datasets can be exploited to achieve high scores without actually detecting any security vulnerabilities.
We conclude that the prevailing problem statement of ML4VD is ill-defined and call into question the internal validity of this growing body of work. Constructively, we call for more effective benchmarking methodologies to evaluate the true capabilities of ML4VD, propose alternative problem statements, and examine broader implications for the evaluation of machine learning and programming analysis research.
@@ I Introduction
In our survey of all ML4VD papers (81) published in the top Software Engineering and Security conferences and journals between January 2020 and December 2024, [^1] Software Engineering conferences: ICSE, FSE, ASE, and ISSTA; Security conferences: S&P, USENIX, NDSS, and CCS; Software Engineering and Security journals: TSE, TOSEM, TDSC, and TIFS. we find that the great majority (88%) of state-of-the-art ML4VD techniques define ML4VD as a function-level binary classification problem: Given only the function $f$, decide whether $f$ contains a security vulnerability or not. The technique with the lowest classification error on the evaluation dataset is considered the best at detecting security vulnerabilities.
@@ 21 TF_LITE_ENSURE_STATUS(context->ResizeTensor(context, output, output_dims));
However, based on our experience, we hypothesized that it might not always be possible to determine whether a function is vulnerable or not without additional context. We call these vulnerabilities context-dependent. Consider the example in Figure 1. If this function from the DiverseVul benchmark dataset [1] is called with num_splits set to zero, it will crash with a division-by-zero in Line 7. However, without knowing whether this function can ever be called with num_splits set to zero, we cannot reliably decide if the division-by-zero could actually be observed. This function parameter might as well be properly validated by every caller function. This context-dependency problem is well-studied for other approaches to vulnerability detection, such as static analysis [2, 3] (where it is handled, e.g., by defining suitable preconditions) or software testing [4] (which distinguishes, e.g., between system- and unit-level testing).
In this paper, we set out to quantify the prevalence of context-dependent vulnerabilities in the most popular datasets and end up revealing a fundamental flaw in the most-widely used evaluation methodology that is underpinning the progress of the nascent research area of ML4VD. We find that the vulnerability of a function cannot be decided without further context for more than 90% of functions. This includes functions with both types of labels, vulnerable or secure: If the right context existed, a function labeled as secure would make the program vulnerable. Respectively, only because the right context exists, a function labeled as vulnerable makes the program actually vulnerable.
Given our findings, we conclude that the prevailing problem statement of ML4VD as a function-level classification problem is inadequate. The reported results in the literature, which are based on this problem statement, may not accurately reflect the true capabilities of the evaluated techniques at the task of vulnerability detection. In other words, there is currently no evidence that ML4VD techniques are actually capable of identifying security vulnerabilities at the function-level.
But why do ML4VD techniques still achieve high scores at this binary classification task when there is demonstrably not enough information in over 90% of samples (even after addressing label inaccuracies)? We identify spuriously correlated features as a potential reason. Training simple models, like a gradient boosting classifier using only word counts and disregarding code structure, we achieved results comparable to those of state-of-the-art ML4VD models. This suggests that ML4VD techniques only appear to perform well due to the chosen evaluation methodology. During classification, ML4VD techniques rely on spuriously correlated features to achieve high scores and do not genuinely detect vulnerabilities.
@@
    We analyze all papers published at the top Software Engineering conferences, Security conferences, and journals over the last five years and find that the great majority (88%) of state-of-the-art ML4VD techniques define ML4VD as a function-level classification problem.
@@
    Why do ML4VD techniques still achieve high scores at the function-level classification problem? We demonstrate that they may rely on spurious features to achieve high scores without genuinely detecting vulnerabilities.
@@ II Background
The context dependency problem—i.e., the determination whether a function causes a program to be vulnerable depends on the context that is external to the given function—has been well-studied in static analysis and software testing.
@@
Static Analysis. We distinguish between inter-procedural and intra-procedural analysis, where the former is concerned with the analysis of the entire system and the latter with the analysis of individual functions. Intra-procedural static analysis tools often struggle with false positives due to context-dependency. This is a core problem in static analysis, which has received much attention in the literature [5]. In practice, the context-dependency check is ultimately delegated to the user of the tool (which, however, similarly complicates the experimental evaluation in the absence of a user). A programmatic approach is to (manually or automatically) annotate the function with a precondition (e.g., in the style of Hoare logic [6]) that encodes assumptions about the valid state space at the start of the function call. Function summaries or type systems provide other mechanisms to encode the valid context under which the function is (meant to be) called. Despite these efforts, discerning the specific conditions under which a bug manifests remains a challenge. For example, a static analyzer might flag potential bugs based on the analysis of code patterns but cannot always discern the specific conditions under which a bug manifests [7, 8]. Hence, Le et al. [2] propose to differentiate between manifest bugs, which are context-independent, and latent bugs, which depend on preconditions in the calling context. Manifest bugs can be reported without making any assumptions about the calling context, whereas latent bugs cannot.
@@
Software Testing. We distinguish between system- and unit-level testing [9], where the former is concerned with testing the entire system [10, 11] and the latter with testing individual system units, like functions [12, 13]. Automated unit test generation often falls short of identifying issues that only occur under specific conditions or in particular environments. As noted by Harrold and Orso [14], unit tests can produce false positives when they are not adequately designed to account for the broader context in which a function operates. This can lead to an inflated number of reported issues that are not actual bugs, thereby complicating the debugging process. Hence, property-based testing [15, 16, 17] requires users to define function preconditions in addition to assertions to ensure that the assertions fire only under valid function parameters (i.e., under valid context).
To summarize, both the static analysis and software testing research communities acknowledge the problem of context-dependency and address it through various techniques aimed at reducing the number of false positives.
Machine Learning (ML4VD). Several colleagues have previously raised concerns about the validity of this growing body of work. Ullah et al. [18] study the performance of LLMs for vulnerability detection and find that results are nondeterminstic, the reasoning is incorrect and unfaithful, and that they perform poorly in real-world scenarios. Risse and Böhme [19] find that the capability of state-of-the-art ML4VD techniques to distinguish between buggy and patched function is barely better than a coin flip. Croft et al. [20] study the quality of benchmark datasets used in the evaluation of ML4VD techniques. They find that a large percentage of vulnerability labels in real-world datasets are in fact inaccurate and many data points were even duplicated (sometimes with inconsistent labels). In this paper, we reproduce their experiment which allows us to identify and study actually vulnerable functions in terms of their context-dependence. However, our question remains: Can the vulnerability of a function be decided without further context in the first place? Only if it can, we need to be concerned about the quality of our datasets and evaluation procedures.
Another stream of works studies the assumption that a vulnerability is localized within a single function. For instance, Sejfia et al. [21] realize that the patches of some vulnerabilities span across multiple “base units” (e.g, functions) and measure the performance of vulnerability detectors to individually classify all units belonging to the same vulnerability as vulnerable—observing a substantial drop in performance. They also recommend to ensure that the train/validation/test split honors the grouping of units by vulnerability. Li et al. [22] study the default assumption that the vulnerability-patching functions are labeled as vulnerable and find that the patch and its ”trigger” (i.e., the code where the vulnerability manifests, e.g., as an array out-of-bounds write, a double-free, etc.) exist in different functions for about a quarter of studied vulnerabilities. [^2] To facilitate this evaluation, they also introduce a slicing-based approach, VulTrigger, to identify the trigger given the patch. If we labeled functions containing the vulnerability-triggering code as vulnerable, function-level vulnerability detectors would turn out much less effective, they find.
However, they still define ML4VD as a unit-level classification problem. How prevalent is the function-level problem statement for ML4VD in the literature (§III)? Is it a valid problem statement to begin with (§IV, §V)? In answering these questions, our paper aims to uncover the shortcomings of current evaluation methodologies and emphasizes the urgent need for more context-aware approaches to accurately assess the true capabilities of ML4VD techniques in detecting security vulnerabilities.
Spurious features. If we find that the decision whether a function is vulnerable or not cannot be made without further context that is external to that function, then how do we explain the high values of classifier performance? Maybe a classifier learns to predict well based on features that are spuriously correlated with the vulnerable label. Indeed, the phenomenon of spurious correlations in ML4VD is well-studied. For instance, Arp et al. [23] provide initial evidence that machine learning techniques can learn to predict labels correctly based on artifacts in code snippets without addressing the actual security task at hand (spurious correlation). Risse et al. [19, 24] demonstrate that ML4VD techniques overfit to label-unrelated features using semantic-preserving transformations and further that making the models robust against semantic-preserving transformation only introduces overfitting to the specific transformations used.
Furthermore, several studies have proposed ML4VD techniques that are aimed to be less affected by spurious correlations. For example, Rahman et al. [25] identify spuriously correlated variable and API names and disable their use during inference time of their model. Cao et al. [26] completely remove code that is not related to the vulnerability from the code of the vulnerable function during training by leveraging patch commit information. They develop a novel graph-based neural network architecture for the function-level vulnerability classification task to classify vulnerabilities based on causal features only.
While existing work studies and minimizes the impact of spurious correlations using structural representations or semantic-preserving transformations of the vulnerable code (i.e., perturbations on train and/or test data), we provide substantial evidence of this phenomenon using an even simpler and technique-independent method: A high measure of classifier performance is achieved even by a simple classifier if there are no program semantics but only word counts available as features. In this way, our work provides an orthoganal explanation of the surprising results of ML4VD even though the problem statement is ill-defined.
@@ III-A Methodology
To ensure a thorough and unbiased analysis of the recent literature on machine learning for vulnerability detection (ML4VD), we adopted a systematic methodology based on established guidelines for conducting literature reviews in Software Engineering research [27, 28]. Our approach encompassed three primary phases: defining the scope, selecting relevant papers, and systematically analyzing the selected papers.
(a) ML4VD papers per problem statement granularity.
(b) ML4VD papers per year.
Figure 2: Literature survey results for the 81 ML4VD papers we identified in the top Software Engineering (SE) and Security conferences and journals. Figure 2a shows how the papers define the problem of ML4VD. Note that a paper may use multiple granularities, which explains why the numbers in Figure 2a do not add up to 100%. Figure 2b shows how many papers were published each year since 2020.
@@
Scope. Our literature survey focuses on the past five years, targeting publications from 2020 to 2024. Given the prominence and impact of certain venues in the fields of Software Engineering and Security, we selected a set of the most impactful conferences and journals. For Software Engineering, we included International Conference on Software Engineering (ICSE), Foundations of Software Engineering (FSE), International Symposium on Software Testing and Analysis (ISSTA), and Automated Software Engineering (ASE). For Security, we incorporated Symposium on Security and Privacy (S&P), Network and Distributed System Security Symposium (NDSS), Computer and Communications Security (CCS), and USENIX Security Symposium (USENIX). Additionally, we considered the top journals in these domains, including Transactions on Software Engineering and Methodology (TOSEM), Transactions on Software Engineering (TSE), Transactions on Information Forensics and Security (TIFS), and Transactions on Dependable and Secure Computing (TDSC). These venues were chosen for their reputation in publishing high-quality and influential research in the Software Engineering and Security fields (CORE [^3] CORE: https://www.core.edu.au/home rank A* or A).
Criteria for Paper Selection. To identify relevant papers, we established clear selection criteria. The papers must be full-length papers published in the proceedings of the listed in-scope conferences or journals between 2020 and 2024. We do not include short papers (e.g. for PhD student competitions), since they usually represent work-in-progress. We also do not include pure literature reviews if they do not present new experimental data. The papers must fit the topic of machine learning for vulnerability detection (ML4VD), which means that they either propose a new ML-based technique for vulnerability detection, or that they present experiments which use existing techniques. We include all types of machine learning techniques, e.g. graph neural networks, large language models, or convolutional neural networks. We explicitly focus on vulnerability detection for software, which excludes related fields, e.g. hardware or blockchain vulnerability detection. To facilitate the paper identification process, we defined a set of keywords that are indicative of the research focus in this area: vulnerability, vulnerable, detect, detection, discovery, machine, learning, artificial, intelligence, AI, ML, deep, graph, neural, network, large, language, model. We scanned the titles and abstracts of all in-scope papers using multiple tools: IEEE Xplore Advanced Search [^4] IEEE Xplore Advanced Search: https://ieeexplore.ieee.org/search/advanced for the IEEE venues (ICSE, ASE, S&P, TSE, TDSC, and TIFS), ACM DL Advanced Search [^5] ACM DL Advanced Search: https://dl.acm.org/search/advanced for the ACM venues (ICSE, FSE, ASE, ISSTA, CCS, and TOSEM) and Google Scholar Advanced Search  [^6] Google Scholar Advanced Search: https://scholar.google.com/ for the independent venues (NDSS and USENIX). This resulted in more than 2500 potential papers, which were flagged for further review.
Filtering Process. As a first step, we manually filtered the flagged titles to specifically exclude those that most likely do not fit the topic of ML4VD. One of the co-authors of this paper, a Software Security researcher with publications at the top Software Engineering and Security conferences on the topic of ML4VD, manually checked each of the paper titles to facilitate the first round of filtering. This resulted in a preliminary list of 109 papers, which we include as supplementary material to this paper. The author then reviewed each of the 109 abstracts and, where necessary, the full texts to assess their relevance based on our selection criteria. This resulted in a final list of 81 papers matching the selection criteria. The 28 excluded papers are either off-topic (e.g. vulnerablity management) or do not leverage ML-based techniques (e.g. static analysis).
@@
To answer research questions 1) and 2), we documented how ML4VD is defined, and the datasets used for empirical evaluation within each of the 81 identified ML4VD papers.
@@ III-B Results
Our literature survey identified 81 papers that met our selection criteria and were included in our analysis. The full list of papers is part of the supplementary material of this paper. 39 papers were published in the Software Engineering conferences, 8 in the Security conferences, and 34 in the journals. The distribution of these papers across the years and conferences, as displayed in Figure 2b, indicates a clear trend: the number of papers focused on ML4VD is increasing annually. This upward trajectory underscores the growing interest and significance of this research area within the Software Engineering and Security communities. Additionally, we found a large variety of machine learning techniques being employed, utilizing different data representations (graph-based, token-based), model architectures (e.g., large language models, graph neural networks, convolutional neural networks), and learning algorithms (e.g., contrastive learning).
Problem Statement. Figure 2a shows how the 81 papers included in our literature survey define the problem of ML4VD. To our surprise, the great majority (88%) of papers across the in-scope Software Engineering conferences (92%), Security conferences (75%) and journals (85%) define ML4VD as a binary function-level classification problem. This problem statement, which focuses on determining whether a given function contains a security vulnerability, has become the dominant problem statement in the field. The yearly trend seems to be increasing since 2021, with 95% of papers relying on a function-level problem statement in 2024 (2020: 100%, 2021: 57%, 2022: 75%, 2023: 88%, 2024: 95%). 15 out of 81 (19%) papers define ML4VD at the line-/statement-level and 13 out of 81 papers (16%) at the inter-procedural slice-level. A inter-procedural slice is a subset of a program’s statements affecting a particular computation, spanning across multiple functions or procedures. Notably, the two most popular problem statements (function-level and line-/statement-level) exclusively rely on information within a single function for classification (intra-procedural). The remaining problem statements (file-, program-, commit-, and repository-level) all rely on inter-procedural information, but are only used by a minority of papers (10 out of 81).
Datasets. Regarding datasets, the majority of the papers relied on a limited set of popular datasets for empirical evaluation. Figure 3 shows all datasets that were used by the 81 papers and the number of times they were used. Specifically, BigVul [29] and Devign [30] were both used by 36 papers, and ReVeal [31] by 22 papers. Notably, 53 out of the 81 papers utilized at least one of these three datasets, reflecting their dominance in the field.
Figure 3: Datasets that are used by ML4VD papers published at the top Software Engineering and Security conferences and journals over the last five years. Datasets that were used only once are displayed as ”Other”.
Literature Survey. The great majority of ML4VD papers (88%) published at top Software Engineering conferences (ICSE, FSE, ISSTA, and ASE), Security conferences (S&P, USENIX, NDSS, and CCS), and journals (TSE, TOSEM, TIFS, and TDSC) over the past five years define ML4VD as a binary classification problem: Given an isolated function $f$, decide whether $f$ contains a security vulnerability. Moreover, three datasets—BigVul, Devign, and ReVeal—are overwhelmingly favored, with 53 out of the 81 papers using at least one of these datasets for empirical evaluation.